OpenStudy (anonymous):

Find the measure of Angle 6. measure of angle 7 = 2x+15 measure of angle 8 = 3x

OpenStudy (anonymous):

please help me i very bad at geometry

OpenStudy (anonymous):

sure, do you have a pic or is this the whole question?

OpenStudy (anonymous):

if there's a pic then could you draw it?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

yes i have to upload the picture it will take a minute

OpenStudy (anonymous):

I'll help then, wait a min

OpenStudy (anonymous):

ok thank you very much

OpenStudy (anonymous):

well, if 7=2x+15, then angle 5 is the same

OpenStudy (anonymous):

im trying to find the measure of angle 8

OpenStudy (anonymous):

and 6=3x so is angle 8

OpenStudy (anonymous):

yeah, but you need to put it all together

OpenStudy (anonymous):

thats the thing i dont understand how to do this im taking this class online

OpenStudy (anonymous):

since a whole circle is 360 degrees, 2x+15+2x+15+3x+3x

OpenStudy (anonymous):

can you simplify that?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

im looking for x right

OpenStudy (anonymous):

simplify it

OpenStudy (anonymous):

and tell me:)

OpenStudy (anonymous):

Look here, what's 3x +3x+2x+2x?

OpenStudy (anonymous):

How many x's?

OpenStudy (anonymous):

3x^2 + 2x^2 right

OpenStudy (anonymous):

so 10x

OpenStudy (anonymous):

yea basically

OpenStudy (anonymous):

so its 10x=30 which x is equal to 3 right

OpenStudy (anonymous):

NO 2 times 15 is 30 so 10x +30= the whole circle, and since a circle has 360 DEGREES

OpenStudy (anonymous):

10x +30=360

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

yes i understand dat so i would subtract 30 from 360 which is 330 and 330 divided by 10

OpenStudy (anonymous):

and what is 330 divided by 10?

OpenStudy (anonymous):

It is 3.3 Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so x=3.3 Are you following along?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Since you're trying to find angle 8, and angle 8 is 3x, just multiply 3.3(x) times 3

OpenStudy (anonymous):

3.3 times 3= 9.9 degrees

OpenStudy (anonymous):

ok but angle is obtuse

OpenStudy (anonymous):

i mean angle 8 is an obtuse angle

OpenStudy (anonymous):

That's true, I'll look at it for a second :)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

x=33

OpenStudy (anonymous):

ok so ill fill x into the equation

OpenStudy (anonymous):

Okay, tell me what you get:)

OpenStudy (anonymous):

If you want I can sum all of it up for you in case you still haven't got the gist:)

OpenStudy (anonymous):

so it would be 3 times 33 which is 99

OpenStudy (anonymous):

You're correct! Do you want me to sum the question up, like how we solved it?

OpenStudy (anonymous):

yes i would like to know how you got 33

OpenStudy (anonymous):

sure, i will

OpenStudy (anonymous):

ok i have another question how did you get the value of x

OpenStudy (anonymous):

First of all, angles 6 and 8 are equal. Angles 5 and 7 are equal too. That's a rule in geometry. If you have four angles formed together like that, then opposite sides are equal

OpenStudy (anonymous):

Then, you look at your 2 equations: 7=2x+15 and 8=3x

OpenStudy (anonymous):

Double both of those exuations and you will get : 10x +30

OpenStudy (anonymous):

Since all of the angles you see added together =360 degrees(because it's a circle) 10x+30+360 degrees

OpenStudy (anonymous):

Can you catch on?

OpenStudy (anonymous):

i got confused when you said double the equations

OpenStudy (anonymous):

okay. lat's get back to there then:)

OpenStudy (anonymous):

Do you get that angles 6 and 8 are equal. Angles 5 and 7 are equal too????????

OpenStudy (anonymous):

ye

OpenStudy (anonymous):

yes i understand that

OpenStudy (anonymous):

So to add up all of the angles, you have to do 2x+15+2x+15+3x+3x Are you catching on?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok i have another problem i dont understand can you please help me

OpenStudy (anonymous):

sure I will help:)

OpenStudy (anonymous):

Go on:)

OpenStudy (anonymous):

Find the measure of Angle 2. measure of angle 1 = x measure of angle 2 = x-6

OpenStudy (anonymous):

angle three is equal to 90 degrees so if im right and this a a supplementary angle then the sum would be 180 right

OpenStudy (anonymous):

First, you need to know that the whole thing is 180 degrees (half of a circle)

OpenStudy (anonymous):

yes i know that the whole thing is 180 and angle 3 is 90 degrees.

OpenStudy (anonymous):

wouldnt i divide 90 by 2 since 90 times 2 is 180 and i need to find what equals 90

OpenStudy (anonymous):

so 90 +2x-6=180 degrees That would be my approach:)

OpenStudy (anonymous):

Now all you have to do is use algebra to solve this equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

90+2x-6=180 =2x=90 =x=45 I just subtracted 90 on both sides:)

OpenStudy (anonymous):

Can you catch on?